Given two events G and H, the probabilities of each occurring are as follows: P(G) = 0.24; P(H) = 0.15; P(H AND G) = 0.01. Using this information:
Find the complement of P (H AND G).

Answer by ikleyn(52894)

.
The complement of P (H AND G) = 1 - P (H AND G) = 1 - 0.01 = 0.99.


All you need to do to solve the problem is to read the condition and identify the terms.